# Symmetry and Topology in the Non-Hermitian Kitaev chain

### v2

@@ I Introduction
In a set of seminal work [71, 61], it was shown that NHSE has a deep connection with topological properties of the system. It was shown that with non-reciprocal hoppings in the Su-Schrieffer-Heeger (SSH) model, the topological phase transition point in a finite open chain does not coincide with that of an infinite (or periodic) chain. The true transition point is governed by non-Bloch bulk-boundary correspondence, which is central to topological non-Hermitian systems. This discovery has inspired a slew of research at the intersection of non-Hermitian physics and topology [60, 20, 49, 75, 74, 52, 53, 10, 36, 72, 42, 67, 11, 31, 60, 37, 70, 45, 14, 77, 29, 39, 40, 66, 23, 44, 33, 32, 46, 3, 34, 16, 13, 47, 57, 48, 28].
This naturally raises the question of whether such behavior is universal in one-dimensional systems. By considering a non-Hermitian Kitaev chain [35, 2, 73, 41, 69, 58, 59, 15, 4] with non-reciprocal nearest neighbor hopping amplitudes ($t_{1}$ and $t_{2}$) and asymmetric superconducting pairing terms ($\Delta_{1}e^{-i\varphi_{1}}$ and $\Delta_{2}e^{i\varphi_{2}}$), see Fig. 1, we show that this is not generally the case. For a non-Hermitian Kitaev chain, cf. Fig. 1, the topological phase transition points for finite open chain coincides with that of an infinite (or periodic) chain. By performing a symmetry analysis, we show that particle-hole symmetry (PHS) always exists across the entire parameter regime and thereby protects the Majorana modes from NHSE. However, we show that particle and hole bulk modes exhibit an equal and opposite skin effect, which is also a consequence of PHS. Finally, we present the phase diagram for the non-Hermitian Kitaev chain in the complex chemical potential ($\mu$) plane.
Figure 1: A caricature of the non-Hermitian Kitaev chain with non-reciprocal left and right hopping amplitudes, $t_{1}$ and $t_{2}$ and asymmetric superconducting pairing terms $\Delta_{1}e^{-i\varphi_{1}}$ and $\Delta_{2}e^{i\varphi_{2}}$. Each site is labeled by a positive integer $j(\in\mathbb{Z}^{+})$ and has a finite on-site (chemical) potential $\mu$.
@@ III.1 Finite non-Hermitian Kitaev chain
$A$ $=\begin{pmatrix}-\mu&0\\ 0&\mu\end{pmatrix}\ \ \,B=\begin{pmatrix}-\frac{t_{1}}{2}&\frac{\Delta_{2}}{2}\\ -\frac{\Delta_{1}}{2}&\frac{t_{2}}{2}\end{pmatrix}$
$C$ $=\begin{pmatrix}-\frac{t_{2}}{2}&-\frac{\Delta_{2}}{2}\\ \frac{\Delta_{1}}{2}&\frac{t_{1}}{2}\end{pmatrix}$, (7)
where we have again assumed for simplicity $\mu,t_{1},t_{2}\in\mathbb{R}$ and $\varphi_{1}=\varphi_{2}=0$. Diagonalizing the Hamiltonian in Eq. (6) reveals a complex band structure, Fig. 2(b)-(c) and the presence of zero-energy modes in certain parameter regime, Fig. 2(a).
@@
Figure 3: Probability density of the zero modes (doubly degenerate ground states) and first excited state in non-Hermitian Kitaev chain of length $L=50$, with open boundary conditions (OBC) in the topological phase i.e., $|\mu|<\frac{1}{2}(t_{1}+t_{2})$. The chemical potential is $\mu=0.2$ and the superconducting pairing terms are set to $\Delta_{1}=2=\Delta_{2}$ with $\varphi_{1}=0=\varphi_{2}$. The zero modes ($\psi_{0}^{(1)}$ and $\psi_{0}^{(2)}$) do not show any skin-effect with non-reciprocal hoppings. However, the first excited state ($\psi_{1}$) shows skin-effect (right or left localization) depending on the hopping parameters. (a) Sum of probability densities of zero mode wave functions, $\psi_{0}^{(1)}$ and $\psi_{0}^{(2)}$ with hopping parameters $t_{1}=1.5$ (left hopping), $t_{2}=1$ (right hopping). (b) Sum of probability densities of zero mode wave functions with reciprocal hoppings $t_{1}=1=t_{2}$. (c) Sum of probability densities of zero mode wave functions with reversed reciprocities in hopping i.e., $t_{1}=1$ and $t_{2}=1.5$. (d) Probability density of the first excited state showing left localization (blue) with $t_{1}(=1.5)>t_{2}(=1)$, right localization (red) with $t_{1}(=1)<t_{2}(=1.5)$ in contrast to the case with reciprocal hoppings $t_{1}=1=t_{2}$ (green).
@@ III.2 Generalized Brillouin zone
The NHSE is an exponential localization of wave functions and is the result of a modified Brillouin zone (BZ) in non-Hermitian systems [71]. Simplest way to obtain the modified Brillouin is to make the substitution $e^{ika}\to\beta$ (note that $a$ is the lattice constant and is set to one), where $|\beta|$ is determined by the non-reciprocities (or the non-Hermiticity) in the hopping parameters [71, 61]. Hamiltonian $H^{\text{NH}}$, Eq. (4) with the substitution $e^{ik}\to\beta$ becomes,
@@
With the modified Hamiltonian, we can solve the eigenvalue equation, $\text{det}(H^{\text{NH}}(\beta)-E)=0$ to obtain the energy eigenvalues as a function of $\beta$. We are interested in understanding the behavior of the zero modes of the modified Hamiltonian, Eq. (8). For $E=0$, we have, $\text{det}H^{\text{NH}}(\beta)=0$ and hence we have the following,
@@
From Eq. (9) we find that if we have $\beta_{0}$ as a solution, then $\beta_{0}^{-1}$ is also a solution. Hence, in general, we have four solutions to Eq. (9): $\beta_{1},\,\beta_{1}^{-1},\,\beta_{2},\,\beta_{2}^{-1}$. Additionally, one can show that the matrix, Eq. (8) can be null only at the boundary of the topological and trivial phase, hence the rank of that matrix is 1 at all points in the topological phase. Hence, corresponding to each $\beta$, there is only one zero mode eigenvector in the topological phase. Thus, the four roots of Eq. (9) will have four zero mode eigenvectors in the topological phase. Specifically, at $\mu=0$, we have $\beta_{2}=-\beta_{1}$. On physical grounds, we expect that the zero mode eigenvectors corresponding to $\beta_{1}(\beta_{1}^{-1})$ will be the same as $\beta_{1}(-\beta_{1}^{-1})$. Consequently, for $E=0$, the two solutions have inverse localization factors $\beta_{1}$ and $\beta_{1}^{-1}$, which means that the solution $\beta_{1}$ ($\beta_{1}^{-1}$) localizes to the right (left) [41, 69, 59]. The total probability density of the zero-energy states is given by $|\psi_{0}^{(1)}|^{2}+|\psi_{0}^{(2)}|^{2}$, where $\psi_{0}^{(1)}$ and $\psi_{0}^{(2)}$ are the normalized zero modes corresponding to factors $\beta_{1}$ and $\beta_{1}^{-1}$ respectively. This reciprocal pairing eliminates the skin effect for zero-energy states.
The existence of the reciprocal solution pair can also be understood by the symmetries of the Kitaev chain. Particle-hole symmetry enforces solutions to be reciprocal pairs [41]. In section III.3, we demonstrate that the particle-hole symmetry exists for arbitrary complex parameters in the Hamiltonian, thus protecting the Majorana edge modes, whenever they exist, for those parameter values. However, for any excited state (bulk mode), the solutions to $\text{det}(H^{\text{NH}}(\beta)-E)=0$ are not in a reciprocal pair. In fact, if $\beta_{1}$ is a solution corresponding to energy $E$, then $\beta_{1}^{-1}$ is a solution corresponding to energy E-E. Thus we expect an equal and opposite non-Hermitian skin effect for the particle and hole wavefunctions, cf. Fig. 4(b).
Figure 4: Probability density of zero mode wavefunctions and bulk mode wavefunctions of the non-Hermitian Kitaev chain with $L=50$, $\mu=0.2$, $t_{1}=1+i$, $t_{2}=0.5+i$, $\Delta_{1}=2+i$ and $\Delta_{2}=1+i$. The zero modes are immune to the NHSE whereas the particle and hole wavefunctions show an equal and opposite skin effect. (a) Sum of probability densities of the zero modes $\psi_{0}^{(1)}$ and $\psi_{0}^{(2)}$ is symmetric and localized at the edges. (b) The dashed red line denotes the probability density of the smallest positive energy (first excited state for a particle i.e., +E>0+E>0) state and the solid blue line is the probability density of the smallest negative energy (first excited state for a hole i.e., E<0-E<0) state.
@@ III.3 Symmetry analysis of non-Hermitian Kitaev chain
In contrast to Hermitian systems, where the condition $H^{\dagger}=H$ (for Hamiltonian) ties together transpose and complex-conjugation operations, non-Hermitian Hamiltonians require that these operations be treated as independent when defining discrete symmetries. For example, when studying time-reversal symmetry (TRS) in a non-Hermitian system, one should address $H^{*}$ and $H^{T}$ separately, corresponding to TRS and its adjoint TRS. Similar statements hold for chiral-symmetry (CS) and particle-hole symmetry (PHS) [33]. Considering arbitrary complex parameters, cf. Eq. (III) and by using Specht’s theorem [26], we find the symmetry classification for the non-Hermitian Kitaev chain as shown in Table 1.
It is important to note that the particle-hole symmetry (PHS) always persists, with arbitrary complex parameters. The Table 1 presented here considers real $\mu$. For complex $\mu$, PHS still holds, and the only other possible symmetries are TRS and CS, see Appendix A. As a result of PHS, the Majorana zero modes exist and are robust across the entire parameter regime, Fig. 3(a)-(c) and Fig. 4(a). However, the bulk modes exhibit NHSE with particle and hole localizing at opposite edges, Fig. 4(b).
@@ III.4 Topological phase diagram for non-Hermitian Kitaev chain
Similar to the Hermitian Kitaev chain [2], we begin by writing our Hamiltonian, Eq. (4) in the Pauli basis i.e., $H(k)=h_{\mathbb{I}}(k)\sigma_{\mathbb{I}}+h_{x}(k)\sigma_{x}+h_{y}(k)\sigma_{y}+h_{z}(k)\sigma_{z}$, where $\sigma_{\mathbb{I}}$ is the identity matrix and $\sigma_{x},\,\sigma_{y},\,\sigma_{z}$ are the Pauli matrices. The coefficients $h_{\mathbb{I}}(k),\,h_{x}(k),\,h_{y}(k)$ and $h_{z}(k)$ are given as follows,
$h_{\mathbb{I}}(k)$ $=i(t_{1}-t_{2})\sin k\,$,
@@
For $t_{1}+t_{2}\geq 2|\mu|$, we have $\nu=-1$, which is identified with the topological phase with two zero modes. In the case of $t_{1}+t_{2}<2|\mu|$, we have $\nu=+1$, which corresponds to the trivial phase, cf. Fig. 2(a)-(c).

### v3

# Symmetry and Topology in a Non-Hermitian Kitaev chain

@@ I Introduction
In a set of seminal work [75, 65], it was shown that NHSE has a deep connection with topological properties of the system. It was shown that with non-reciprocal hoppings in the Su-Schrieffer-Heeger (SSH) model, the topological phase transition point in a finite open chain does not coincide with that of an infinite (or periodic) chain. The true transition point is governed by non-Bloch bulk-boundary correspondence, which is central to topological non-Hermitian systems. This discovery has inspired a slew of research at the intersection of non-Hermitian physics and topology [64, 21, 52, 79, 78, 55, 56, 10, 37, 76, 43, 71, 11, 32, 64, 38, 74, 48, 14, 81, 30, 40, 41, 70, 24, 45, 34, 33, 49, 3, 35, 17, 13, 50, 61, 51, 29, 16].
Figure 1: A caricature of a non-Hermitian Kitaev chain with non-reciprocal left and right hopping amplitudes, $t_{1}$ and $t_{2}$, and asymmetric superconducting pairing terms, $\Delta_{1}e^{-i\varphi_{1}}$ and $\Delta_{2}e^{i\varphi_{2}}$. The right hopping is represented by a dotted arrow and the left hopping is represented by a solid arrow. Similarly, asymmetry in the superconducting pairing terms is represented by solid and dotted wave-like arrows. Each site is labeled by an integer $j$ and has a finite on-site (chemical) potential $\mu$.
This naturally raises the question of whether such behavior is universal in one-dimensional systems. By considering a non-Hermitian Kitaev chain [36, 2, 77, 42, 73, 62, 63, 15, 4, 47] with non-reciprocal nearest neighbor hopping amplitudes ($t_{1}$ and $t_{2}$) and asymmetric superconducting pairing terms ($\Delta_{1}e^{-i\varphi_{1}}$ and $\Delta_{2}e^{i\varphi_{2}}$), see Fig. 1, we show that this is not generally the case. For a non-Hermitian Kitaev chain, cf. Fig. 1, the topological phase transition points for finite open chain coincides with that of an infinite (or periodic) chain. By performing a symmetry analysis, we show that particle-hole symmetry (PHS) always exists across the entire parameter regime and thereby protects the Majorana modes from NHSE. However, we show that particle and hole bulk modes exhibit an equal and opposite skin effect, which is also a consequence of PHS. Finally, we present the phase diagram for the non-Hermitian Kitaev chain in the complex chemical potential ($\mu$) plane.
@@ III.1 Finite non-Hermitian Kitaev chain
$A=\begin{pmatrix}-\mu&0\\ 0&\mu\end{pmatrix}\,\,B=\begin{pmatrix}-\frac{t_{1}}{2}&\frac{\Delta_{2}}{2}\\ -\frac{\Delta_{1}}{2}&\frac{t_{2}}{2}\end{pmatrix}\,\,C=\begin{pmatrix}-\frac{t_{2}}{2}&-\frac{\Delta_{2}}{2}\\ \frac{\Delta_{1}}{2}&\frac{t_{1}}{2}\end{pmatrix}$,
where we have again assumed for simplicity $\mu,t_{1},t_{2}\in\mathbb{R}$ and $\varphi_{1}=\varphi_{2}=0$. Diagonalizing the Hamiltonian in Eq. (A1) reveals a complex band structure, Fig. 2(b)-(c) and the presence of zero-energy modes in certain parameter regime, Fig. 2(a).
@@
Figure 3: Probability density of the zero modes (doubly degenerate ground states) and first excited state in non-Hermitian Kitaev chain of length $L=50$, with open boundary conditions (OBC) in the topological phase i.e., $|\mu|<\frac{1}{2}(t_{1}+t_{2})$. The chemical potential is $\mu=0.2$ and the superconducting pairing terms are set to $\Delta_{1}=2=\Delta_{2}$ with $\varphi_{1}=0=\varphi_{2}$. The zero modes ($\psi_{0}^{(1)}$ and $\psi_{0}^{(2)}$) do not show any skin-effect with non-reciprocal hoppings. However, the first excited state ($\psi_{1}^{+}$) shows skin-effect (right or left localization) depending on the hopping parameters. (a) Sum of probability densities of zero mode wave functions, $\psi_{0}^{(1)}$ and $\psi_{0}^{(2)}$ with hopping parameters $t_{1}=1.5$ (left hopping), $t_{2}=1$ (right hopping). (b) Sum of probability densities of zero mode wave functions with reciprocal hoppings $t_{1}=1=t_{2}$. (c) Sum of probability densities of zero mode wave functions with reversed reciprocities in hopping i.e., $t_{1}=1$ and $t_{2}=1.5$. (d) Probability density of the first excited state $\psi_{1}^{+}$ with smallest non-vanishing energy i.e., $|E|\neq 0$ and $\text{Re}(E)>0$, showing left localization (blue) with $t_{1}(=1.5)>t_{2}(=1)$, right localization (red) with $t_{1}(=1)<t_{2}(=1.5)$ in contrast to the case with reciprocal hoppings $t_{1}=1=t_{2}$ (green).
@@ III.2 Generalized Brillouin zone
The NHSE is an exponential localization of wave functions and is the result of a modified Brillouin zone (BZ) in non-Hermitian systems [75]. The simplest way to obtain the modified Brillouin is to make the substitution $e^{ika}\to\beta$ (note that $a$ is the lattice constant and is set to one), where $|\beta|$ is determined by the non-reciprocities (or the non-Hermiticity) in the hopping parameters [75, 65]. Hamiltonian $H^{\text{NH}}$, Eq. (4) with the substitution $e^{ik}\to\beta$ becomes,
@@
With the modified Hamiltonian Eq. (7), we can solve the eigenvalue equation, $\text{det}(H^{\text{NH}}(\beta)-E)=0$ to obtain the energy eigenvalues as a function of $\beta$. We are interested in understanding the behavior of the zero modes of the modified Hamiltonian, Eq. (7). For $E=0$, we have, $\text{det}H^{\text{NH}}(\beta)=0$ and hence we have the following,
@@
From Eq. (8) we find that if we have $\beta_{0}$ as a solution, then $\beta_{0}^{-1}$ is also a solution. Hence, in general, we have four solutions to Eq. (8): $\beta_{1},\,\beta_{1}^{-1},\,\beta_{2},\,\beta_{2}^{-1}$. Additionally, one can show that the matrix, Eq. (7) can be null only at the boundary of the topological and trivial phase (this can be seen directly from Eq. (7), which becomes null only at $\beta=\pm 1$ and $\mu=\mp\frac{t_{1}+t_{2}}{2}$), hence the rank of the matrix in Eq. (7) is 1 at all points within the topological phase. Hence, corresponding to each $\beta$, there is only one zero mode eigenvector in the topological phase. Thus, the four roots of Eq. (8) will have four zero mode eigenvectors in the topological phase. Specifically, at $\mu=0$, we have $\beta_{2}=-\beta_{1}$. On physical grounds, we expect that the zero mode eigenvectors corresponding to $\beta_{1}(\beta_{1}^{-1})$ will be the same as $\beta_{1}(-\beta_{1}^{-1})$. Consequently, for $E=0$, the two solutions have inverse localization factors $\beta_{1}$ and $\beta_{1}^{-1}$, which means that the solution $\beta_{1}$ ($\beta_{1}^{-1}$) localizes to the right (left) [42, 73, 63]. The total probability density of the zero-energy states is given by $|\psi_{0}^{(1)}|^{2}+|\psi_{0}^{(2)}|^{2}$, where $\psi_{0}^{(1)}$ and $\psi_{0}^{(2)}$ are the normalized zero modes corresponding to factors $\beta_{1}$ and $\beta_{1}^{-1}$ respectively. This reciprocal pairing eliminates the skin effect for zero-energy states, as seen in Fig. 3(a)-(c). The reciprocal solution pair for Eq. (8) is a result of particle-hole symmetry in the non-Hermitian Kitaev chain [42]. However, for any excited state or bulk mode, the solutions to $\text{det}(H^{\text{NH}}(\beta)-E)=0$ are not in a reciprocal pair. In fact, if $\beta_{1}$ is a solution corresponding to energy $E$, then $\beta_{1}^{-1}$ is a solution corresponding to energy E-E. Hence, eigenstates with energies $\pm E$ exhibit equal and opposite non-Hermitian skin effect, as shown in Fig. 4, and also see appendix A for more details.
The solutions for $\beta$ in Eq. (8) encode the localization length of the Majorana zero modes. This can be understood in a simple setting by considering $\mu=0$ and setting $\varphi_{1}=\varphi_{2}$. As a result Eq. (8) reduces to a quadratic equation for $\beta^{2}$ with reciprocal roots,
$\beta^{2}_{\pm}=\frac{t_{1}^{2}+t_{2}^{2}+2\Delta_{1}\Delta_{2}\pm\sqrt{(t_{1}^{2}-t_{2}^{2})^{2}+4(t_{1}+t_{2})^{2}\Delta_{1}\Delta_{2}}}{2(\Delta_{1}\Delta_{2}-t_{1}t_{2})}.$ (10)
In non-Hermitian lattices with skin effect, the wavefunctions go as $\psi_{j}\sim\beta^{j}$ [75], and the localization length is given by $a/\big|\log|\beta|\big|$, where $a$ is the lattice constant which is set to 1 for simplicity. Thus, the localization length for the two Majorana zero modes are $1/\big|\log|\beta_{+}|\big|$ and $1/\big|\log|\beta_{-}|\big|$, which are equal since we have $\beta_{+}\beta_{-}=1$. Thus, the zero modes are always exponentially localized at the two edges of the chain. In the Hermitian case where $t_{1}=t_{2}=t$ and $\Delta_{1}=\Delta_{2}=\Delta$ in Eq. (10) we find $1/\big|\log|(\Delta+t)/(\Delta-t)|\big|$ as the localization length for the Majorana modes as expected from previous analysis [2].
Next, we discuss the role of $\Delta_{1}$ and $\Delta_{2}$ in the physics of the non-Hermitian Kitaev chain, especially in the robustness of the Majorana modes. One might expect that setting $\Delta_{1}$ and $\Delta_{2}$ arbitrarily small reduces the non-Hermitian Kitaev chain to the Hatano-Nelson model, where all states exhibit the skin effect. However note that the role of even arbitrarily small superconducting pairing terms in the Hamiltonian Eq. (1) and Eq. (III) is to open a gap in the spectrum. The closing and reopening of this bulk gap drives the system from the topological phase (with Majorana zero modes at the edges) to the trivial phase (without zero modes), and vice-versa. As discussed previously, Majorana modes are exponentially localized at the edges of the chain as long as $\Delta_{1,2}\neq 0$. In contrast, the Hatano-Nelson model does not have a topological phase transition.
The Hamiltonian of the non-Hermtian Kitaev chain with $\Delta_{1,2}\neq 0$ takes a Bogoliubov-de-Gennes (BdG) form, Eq. (4), which is another manifestation of the particle-hole symmetry in the system. The BdG form holds for arbitrary complex parameters, which implies an unbroken paricle-hole symmetry throughout the entire complex parameter regime, as we shall demonstrate explicitly in the following subsection III.3. Hence, we see that $\Delta_{1,2}\neq 0$ makes the system very different from the Hatano-Nelson model, and leads to significantly different physics. As a result of particle-hole symmetry, we get reciprocal roots for $\beta$, leading to an equal and opposite skin effect for particles at energy +E+E and holes at energy E-E. As Majorana zero modes are an equal superposition of particles and holes at zero energy, the skin effect is neutralized for these modes, and we get symmetrically localized Majorana zero modes at the edges, cf. Fig. 3(a)-(c).
@@ III.3 Symmetry analysis of non-Hermitian Kitaev chain
In contrast to Hermitian systems, where the condition $H^{\dagger}=H$ (for Hamiltonian) ties together transpose and complex-conjugation operations, non-Hermitian Hamiltonians require that these operations be treated as independent when defining discrete symmetries. For example, when studying time-reversal symmetry (TRS) in a non-Hermitian system, one should address $H^{*}$ and $H^{T}$ separately, corresponding to TRS and its adjoint TRS. Similar statements hold for chiral-symmetry (CS) and particle-hole symmetry (PHS) [34]. Considering arbitrary complex parameters, cf. Eq. (III) and by using Specht’s theorem [27], we summarize the symmetries of the non-Hermitian Kitaev chain in Table 1.
It is important to note that the particle-hole symmetry (PHS) always persists, with arbitrary complex parameters. The Table 1 presented here considers real $\mu$. For complex $\mu$, PHS still holds, and the only other possible symmetries are TRS and CS, see Appendix B. As a result of PHS, the Majorana zero modes exist and are robust across the entire parameter regime, Fig. 3(a)-(c) and Fig. 4(a). However, the bulk modes exhibit NHSE with wavefunctions of opposite energies localizing at opposite edges, Fig. 4(b). To see this, note that PHS on the energy spectrum implies that $E(k)$ and E(k)-E(-k) come in pairs. Once we promote $e^{ik}$ to $\beta$, this implies that $E(\beta)$ and $E(\beta^{-1})$ come in pairs, i.e., the states with energy +E+E and E-E localize at opposite ends. We emphasize that Fig. 4 (and Fig. A1 in Appendix A) demonstrates that the symmetric localization of the Majorana zero modes and the equal and opposite skin effect for eigenstates with energy $\pm E$ for arbitrarily complex parameters in comparison to Fig. 3, where we had $\Delta_{1}=\Delta_{2}$, and all real parameters.
Figure 4: Probability density of zero mode wavefunctions and eigenstates with smallest non-vanishing energy ($\pm|E|\neq 0$), of the non-Hermitian Kitaev chain with $L=50$, $\mu=0.2$, $t_{1}=1+i$, $t_{2}=0.5+i$, $\Delta_{1}=2+i$ and $\Delta_{2}=1+i$. The zero modes are immune to the NHSE whereas the nonzero energy wavefunctions show an equal and opposite skin effect. (a) Sum of probability densities of the zero modes $\psi_{0}^{(1)}$ and $\psi_{0}^{(2)}$ is symmetric and localized at the edges. (b) $\psi_{1}^{+}$ and $\psi_{1}^{-}$ are the energy eigenstates corresponding to complex energies $\pm E$, $\text{Re}(E)>0$ such that $|E|$ is the smallest non-zero absolute value of energy in the spectrum i.e., $|E|\neq 0$. The solid blue line and the dashed red lines denote the probability densities of $\psi_{1}^{+}$ and $\psi_{1}^{-}$ respectively and exhibit opposite skin effect.
@@ III.4 Topological phase diagram for non-Hermitian Kitaev chain
Similar to the Hermitian Kitaev chain [2], we begin by writing our Hamiltonian, Eq. (4) in the Pauli basis i.e., $H(k)=h_{\mathbbm{1}}(k)\mathbbm{1}+h_{x}(k)\sigma_{x}+h_{y}(k)\sigma_{y}+h_{z}(k)\sigma_{z}$, where $\mathbbm{1}$ is the identity matrix and $\sigma_{x},\,\sigma_{y},\,\sigma_{z}$ are the Pauli matrices. The coefficients $h_{\mathbbm{1}}(k),\,h_{x}(k),\,h_{y}(k)$ and $h_{z}(k)$ are given as follows,
$h_{\mathbbm{1}}(k)$ $=i(t_{1}-t_{2})\sin k\,$,
@@
For $t_{1}+t_{2}\geq 2|\mu|$, we have $\nu=-1$, which is identified with the topological phase with two zero modes. In the case of $t_{1}+t_{2}<2|\mu|$, we have $\nu=+1$, which corresponds to the trivial phase, cf. Fig. 2(a)-(c). In the most general parameter regime, $t_{1}\neq t_{2},t_{1}\neq t_{2}^{*},\varphi_{1}\neq\varphi_{2}$, PHS is the only symmetry of the non-Hermitian Kitaev chain (refer Table 1). Using the symmetry-based topological classification of Ref. [35], we observe that the non-Hermitian Kitaev chain is in class D, and hence has a $\mathbb{Z}_{2}$ topological invariant, see Appendix B for more details. It is also worth noting that in experimental setups, the physical observable to detect Majorana zero modes appearing at the ends of a superconducting wire is fermion parity [1], which also corresponds to a $\mathbb{Z}_{2}$ topological invariant.